# BayesFLo: Bayesian fault localization of complex software systems

### v2

@@ 4.1 An Alternate Formulation
$E_{(\mathbf{i},\mathbf{j})}=\{\textup{for each $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m}\setminus\mathcal{C}_{{\rm TP}}$ such that $Z_{c}=1$}\}.$ (9)
@@
The proof of this proposition can be found in Appendix A. There are two key advantages of the alternate form (10) over the brute-force approach (7). First, its numerator can be directly computed via Equation (2) with little work. Second, its denominator $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ can be effectively computed via a novel connection to a related minimal set covering problem for bipartite graphs (Asratian et al.,, 1998), which we show below.
@@ 4.2 Enumerating Minimal Covers
$\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})=\mathbb{P}(\{Z_{c}=1\textup{ for all }c\in\tilde{\mathcal{C}}\}\textup{, for at least one minimal cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$}).$ (11)
@@
To compute (11), a natural approach is to first enumerate all minimal covers of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$. Fortunately, the set cover problem for bipartite graphs has been well-studied in the literature, and efficient polynomial-time algorithms have been developed for finding minimal covers (Skiena,, 1998; Hopcroft and Karp,, 1973). Leveraging such developments can thus greatly speed up the brute-force approach for posterior probability computation (see Equation (7)), which is doubly-exponential in complexity and thus infeasible for even small software systems. With recent developments in integer programming algorithms (Wolsey,, 2020), a popular strategy for finding minimal set covers is to formulate and solve this problem as an integer linear program (ILP; Schrijver,, 1998). We adopt such a strategy below.
@@
$\begin{split}\operatorname*{arg\,max}\;1&\text{\quad s.t. \quad}\\ &z_{c}\in\{0,1\}\;\text{for all }c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},\quad l_{g,m}\in\{0,1\}\;\text{for all }g\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})},\\ \text{[C1] }&\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}}z_{c}\cdot\mathbb{I}(m\in\mathcal{M}_{c})\geq 1\;\text{for all }m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})},\\ \text{[C2] }&\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},c\neq g}z_{c}\cdot\mathbb{I}(m\in\mathcal{M}_{c})\leq|\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}|(1-l_{g,m})\;\text{for all }g\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})},\\ \text{[C3] }&\sum_{m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}}l_{g,m}\geq 1\;\text{for all }g\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}.\end{split}$
@@
One appealing property of the integer feasible program (12) is that the objective is (trivially) linear and all constraints are linear in the binary decision variables. Such an integer linear program thus admits nice structure for efficient large-scale optimization, particularly via recent developments in cutting plane and branch-and-bound algorithms (Balas et al.,, 1993; Stidsen et al.,, 2014). In our later implementation, we made use of the GurobiPy package in Python (Gurobi Optimization, LLC,, 2023), which implements state-of-the-art optimization solvers for large-scale integer programming. Gurobi is widely used for solving large-scale scheduling problems in the industry, including for the National Football League (North,, 2020) and Air France (Richard,, 2020). Here, with the ILP formulation (12), Gurobi can solve for a feasible minimal set cover in minutes for our later case studies. This formulation thus provides an efficient strategy for computing the desired probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$.
@@
where $\oplus$ is the XOR operator. This new constraint [C4] ensures the next cover is distinct from the previous found cover. To see why, note that $d_{c}$ equals 1 only if the binary variables $z_{c}$ and $\tilde{z}_{c}$ are different; the inequality constraint in [C4] thus ensures all considered covers are different from the previous cover $\{\tilde{z}_{c}\}_{c\in\mathcal{C}_{\rm TF,(\mathbf{i},\mathbf{j})}}$. The resulting problem is still an integer linear program, as XOR can naturally be expressed as linear constraints (Magee and Glover,, 1996). More specifically, the XOR condition in [C4] can be equivalently expressed as:
$\begin{split}d_{c}\geq z_{c}-\tilde{z}_{c},\quad d_{c}\geq\tilde{z}_{c}-z_{c},\quad d_{c}\leq z_{c}+\tilde{z}_{c},\quad d_{c}\leq 2-z_{c}-\tilde{z}_{c},\quad c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},\end{split}$ (13)
which is clearly linear in the binary decision variables $\{z_{c}\}_{c\in\mathcal{C}_{\rm TF,(\mathbf{i},\mathbf{j})}}$.
With this, if a feasible solution is found for the ILP (12) with [C4], the optimization solver will return a distinct minimal cover, which we add to the collection. If not, the solver will instead return a “dual certificate” (Güzelsoy et al.,, 2010) that guarantees the ILP has no feasible solutions; such a certificate is possible via the linear nature of the integer program. One then iteratively solves the ILP (12) with constraint [C4] (modified to exclude two or more found covers) until the solver returns a dual certificate, in which case no feasible solutions are possible and thus all minimal covers have been enumerated.
@@ 4.3 Computing Root Cause Probabilities
$\begin{split}\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})&=\mathbb{P}(\{Z_{c}=1\textup{ for all }c\in\tilde{\mathcal{C}}\}\textup{, for at least one $\tilde{\mathcal{C}}\in\mathcal{V}$})\\ &=\sum_{\text{cover }\tilde{\mathcal{C}}\in\mathcal{V}}\;\prod_{c\in\tilde{\mathcal{C}}}p_{c}-\sum_{\text{covers }\tilde{\mathcal{C}},\tilde{\mathcal{C}}^{\prime}\in\mathcal{V}}\;\prod_{c\in\tilde{\mathcal{C}}\cup\tilde{\mathcal{C}}^{\prime}}p_{c}+\cdots+(-1)^{|\mathcal{V}|}\prod_{c\in{\tilde{\mathcal{C}}_{1}\cup\cdots\cup\tilde{\mathcal{C}}_{|\mathcal{V}|}}}p_{c},\end{split}$
@@
which bypasses the need for computing higher-order terms involving more than two covers. Note that, by the inclusion-exclusion principle, the right-hand side of (15) underestimates the probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$. This is by design: from (10), this then results in a slight overestimation of the posterior root cause probability $\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})$. From a risk perspective, this is more preferable than an approximation procedure that underestimates such probabilities.
@@ 4.4 Algorithm Summary
Next, we compute posterior root cause probabilities within each category. For TP combinations, this is trivially zero as such combinations were cleared in passed cases. For UT combinations, this can be set as the prior probabilities from (2), as no information can be gleaned on such combinations from the test data. For TF combinations, its posterior probabilities can be computed via the minimal set cover approach in Section 4. Finally, with posterior probabilities computed, we can then rank the potential root causes in terms of their probabilities, which can be used for guiding software diagnosis; more on this later. Figure 3 (right) visualizes the full workflow for the BayesFLo procedure.
We conclude this section with a discussion of computation time for BayesFLo. When the test data consists of only a single failed test run, the BayesFLo procedure from Algorithm 1 runs very quickly in our experience. For our later Case Study 2, where only one failed test run is observed (see Table 5), all posterior BayesFLo probabilities can be computed in several seconds, which is comparable in time to the state-of-the-art JMP Covering Analysis. When the test data consists of multiple failed runs, BayesFLo can require considerably more computation time, particularly for minimal cover enumeration. For our later Case Study 1, where multiple failed runs are observed (see Table 2), posterior BayesFLo probabilities require several hours of computation, whereas the JMP Covering Analysis takes only seconds.
This increased computation time for BayesFLo in the case of multiple failed test runs is, however, not a major bottleneck for three reasons. First, the key advantage of BayesFLo over the state-of-the-art lies in its improved probabilistic analysis, which permits confident fault localization with a small number of costly test runs. Thus, while our analysis may require more computation, this is in many cases preferable to committing tens of thousands of dollars on investigating many tied suspicious combinations, as shown in later case studies. Second, as alluded to earlier, we presume that at time of performing BayesFLo, the system has been sufficiently tested such that few failed runs are expected. Finally, with the increasing availability of multi-core processors, the computation for BayesFLo can be considerably accelerated by exploiting such parallelism within the GurobiPy package; see Glockner, (2015) for details.
@@
5:For each TF combination $(\mathbf{i},\mathbf{j})$, enumerate minimal covers for the failed cases in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, then compute its posterior root cause probability using Equation (15).
@@ 5 Case Studies
We now explore the effectiveness of BayesFLo for tackling our motivating fault localization case study on the Traffic Alert and Collision Avoidance System (Section 2), as well as another case study on the JMP Easy DOE platform (Jones and Sall,, 2011). The first case study illustrates the importance of a flexible Bayesian modeling framework that integrates structural knowledge for disentangling the many tied combinations from existing analyses. The second explores the importance of integrating domain knowledge to disentangle tied combinations for confident fault localization.
@@ 5.1 Case Study 1: Traffic Alert and Collision Avoidance System
Consider first the motivating case study in Section 2, on the fault localization of the Traffic alert and Collision Avoidance System (TCAS). As mentioned earlier, TCAS targets the avoidance of potential mid-air collisions between aircraft, and its reliable performance is essential for mitigating catastrophic disasters (see, e.g., Davies and Yousif,, 2025). In what follows, we consider a specific program – the Siemens test suite in Griesmayer et al., (2007) – for fault localization. Its $I=12$ input parameters with corresponding levels are summarized in Table 1. We consider the “TCAS v2” set-up from Griesmayer et al., (2007), which contains a fully-documented fault that results in an incorrect resolution advisory; if unaddressed, such a fault may lead to potential aircraft collisions. Here, it is known (Griesmayer et al.,, 2007) that such a failure arises from a single root cause, namely, the three-factor combination of factors 12, 8 and 9:
[…]
With this set-up, we then construct a strength-2 covering array with the input factors and levels from Table 1, using a run size of $M=17$. Table 2 shows the corresponding test outcomes for each test run. Figure 2 outlines the fault localization analysis from the JMP Covering Array module, which we treat as the state-of-the-art approach that integrates several existing methods. Recall that such an analysis has several limitations. First, its top ten combinations are all tied in terms of suspiciousness (i.e., its failure count); the budget needed to investigate all ten combinations can thus be prohibitively expensive, as the debugging of each combination can require thousands of dollars (Bush,, 1990). Second, this analysis is deterministic and offers no insight on the probability that a suspicious combination is indeed a root cause given observed test outcomes. Such probabilistic analysis facilitates confident fault localization and permits the disentangling of tied combinations for timely diagnosis.
To implement BayesFLo, we first need a careful specification of its prior distribution in (2). Here, we have little prior domain knowledge besides the belief that root causes are rare. To reflect this, we adopt the weakly-informative prior of $p_{(i,j)}=1/30$, where the denominator $\sum_{i=1}^{I}J_{i}=30$ is the total number of levels for all $I=12$ factors. In what follows, we consider root cause combinations with at most three factors; this not only reduces computation, but is also guided by prior observations that root causes are typically of lower order for such systems (Kuhn et al.,, 2004). Posterior root cause probabilities are then computed via Algorithm 1.
Top Suspicious Combinations (BayesFLo)
Factors Levels Posterior Probability
12, 8, 9 1, 399, 640 0.55
8, 7 399, 1 0.20
9, 7 640, 1 0.20
7, 10 1, 0 0.20
9, 1 640, 300 0.10
9, 11 640, 2 0.10
7, 10 0, 1 0.10
7, 2 0, 0 0.10
7, 6 0, 2 0.10
1, 10 300, 1 0.10
10, 11 1, 2 0.10
@@
Table 3: Top suspicious combinations from the BayesFLo analysis in the TCAS case study, ranked by decreasing posterior root cause probabilities. Remaining combinations all have posterior probabilities less than $7.0\times 10^{-3}$. Here, the true root cause is bolded.
Table 3 shows the top suspicious combinations from the BayesFLo analysis, ranked by its posterior root cause probabilities. There are two key observations to note. First, compared to the JMP analysis (Figure 2), BayesFLo returns a more informative list that disentangles tied combinations: it identifies the three-factor combination (bolded in Table 3) as the most suspicious with a posterior probability of 0.55, followed by three two-factor combinations with posterior probabilities of 0.20. This top three-factor combination from BayesFLo is precisely the true root cause (16); test engineers can thus quickly identify the true root cause from our analysis, which greatly reduces costs. This highlights the importance of a flexible probabilistic modeling framework that embeds structural principles for disentangling many tied combinations. Second, BayesFLo provides a probabilistic quantification of combination suspiciousness via its posterior root cause probabilities, including for higher-order combinations that would not have been considered by existing methods. Here, beyond the top 11 combinations in Table 3, the remaining all have posterior probabilities less than $7.0\times 10^{-3}$. Such combinations can thus be safely ignored for debugging, which is indeed the correct decision here. This notion of confidence is not provided by existing methods, which may cause greatly delayed root cause diagnosis, as discussed below.
To contrast, the JMP analysis (Figure 2) yields a much more muddled picture for fault localization. In addition to the top ten tied combinations (which are costly to debug), note that such combinations do not include the true root cause (16), which is a three-factor combination. Thus, even if a test engineer invests resources to debug the top ten combinations, such resources would be wasted and additional resources are needed to identify the true root cause. Furthermore, with this analysis, it is unclear when the tester can confidently stop the debugging of suspicious combinations, as it does not come with a probabilistic quantification of uncertainty. These issues can thus result in costly and delayed fault localization of TCAS; BayesFLo rectifies such issues via a carefully-specified Bayesian modeling framework.
5.2 Case Study 2: JMP Easy DOE Platform

### v3

@@ 4.1 An Alternate Formulation
[…]
@@
The proof of this proposition can be found in Appendix A. There are two key advantages of the alternate form (10) over the brute-force approach (7). First, its numerator can be directly computed via Equation (2) with little work. Second, its denominator $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ can be efficiently computed via a novel connection to a related minimal set covering problem for bipartite graphs (Asratian et al., 1998), which we show below.
@@ 4.2 Enumerating Minimal Covers
[…]
@@
To compute (11), a natural approach is to first enumerate all minimal covers of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$. Fortunately, the set cover problem for bipartite graphs has been well-studied in the literature, and efficient polynomial-time algorithms have been developed for finding minimal covers (Skiena, 1998; Hopcroft and Karp, 1973). Leveraging such developments can thus greatly speed up the brute-force approach for posterior probability computation (see Equation (7)), which is doubly-exponential in complexity and thus infeasible for even small software systems. With recent developments in integer programming algorithms (Wolsey, 2020), a popular strategy for finding minimal set covers is to formulate and solve this problem as an integer linear program (ILP; Schrijver, 1998). We adopt such a strategy below.
@@
[…]
@@
One appealing property of the integer feasible program (12) is that the objective is (trivially) linear and all constraints are linear in the binary decision variables. Such an integer linear program thus admits nice structure for efficient large-scale optimization, particularly via recent developments in cutting plane and branch-and-bound algorithms (Balas et al., 1993; Stidsen et al., 2014). In our later implementation, we made use of the GurobiPy package in Python (Gurobi Optimization, LLC, 2023), which implements state-of-the-art optimization solvers for large-scale integer programming. Gurobi is widely used for solving large-scale scheduling problems in the industry, including for the National Football League (North, 2020) and Air France (Richard, 2020). Here, with the ILP formulation (12), Gurobi can solve for a feasible minimal set cover in seconds for our later case studies. This formulation thus provides an efficient strategy for computing the desired probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$.
@@
where $\oplus$ is the XOR operator. This new constraint [C4] ensures the next cover is distinct from the previous found cover. To see why, note that $d_{c}$ equals 1 only if the binary variables $z_{c}$ and $\tilde{z}_{c}$ are different; the inequality constraint in [C4] thus ensures all considered covers are different from the previous cover $\{\tilde{z}_{c}\}_{c\in\mathcal{C}_{\rm TF,(\mathbf{i},\mathbf{j})}}$. The resulting problem is still an integer linear program, as XOR can naturally be expressed as linear constraints (Magee and Glover, 1996). More specifically, the XOR condition in [C4] can be equivalently expressed as:
$\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}}(1-\tilde{z}_{c})\circ z_{c}\geq 1,\quad\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}}\tilde{z}_{c}-\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}}\tilde{z}_{c}\circ z_{c}\geq 1,$ (13)
where $\circ$ is the element-wise product. This is linear in the binary variables $\{z_{c}\}_{c\in\mathcal{C}_{\rm TF,(\mathbf{i},\mathbf{j})}}$.
With this, if a feasible solution is found for the ILP (12) with [C4], the optimization solver will return a distinct minimal cover, which we add to the collection. If not, the solver will instead return a “dual certificate” (Güzelsoy et al., 2010) that guarantees the ILP has no feasible solutions; such a certificate is possible via the linear nature of the integer program. One then iteratively solves the ILP (12) with constraint [C4] (modified to exclude two or more found covers) until the solver returns a dual certificate, in which case no feasible solutions are possible and thus all minimal covers have been enumerated.
@@ 4.3 Computing Root Cause Probabilities
[…]
@@
which bypasses the need for computing higher-order terms involving more than two covers. Note that, by the inclusion-exclusion principle, the right-hand side of (15) underestimates the probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$. This is by design: from (10), this then results in a slight overestimation of the posterior root cause probability $\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})$. From a risk perspective, this is preferable to an approximation procedure that underestimates such probabilities.
@@ 4.4 Algorithm Summary
Next, we compute posterior root cause probabilities within each category. For TP combinations, this is trivially zero as such combinations were cleared in passed cases. For UT combinations, this can be set as the prior probabilities from (2), as no information can be gleaned on such combinations from the test data. For TF combinations, their posterior probabilities can be computed via the minimal set cover approach in Section 4. Finally, we can then rank the potential root causes in terms of their posterior probabilities, which can be used for planning downstream debugging efforts; more on this later.
While it is difficult to establish the theoretical complexity of our algorithm, our implementation in later case studies suggests that it is quite efficient in practice. For both case studies (the first with $I=12$ factors and two failed test runs, and the second with $I=15$ factors and six failed test runs), the desired BayesFLo posterior probabilities can be computed within a minute on a single-core desktop processor. This shows that, with a careful integration of recent integer programming and graph representation tools, one can efficiently perform Bayesian inference for this challenging high-dimensional fault localization problem. With the availability of multi-core processors, BayesFLo can be further accelerated by exploiting such parallelism within the GurobiPy package; see Glockner (2015) for details.
Finally, we discuss how the posterior probabilities from BayesFLo can help guide actionable decisions for confident fault localization. One important use is for planning the subsequent debugging effort for test engineers. Let $\alpha$ be a user-specified significance level, where suspicious combinations with posterior probabilities below $\alpha$ can be confidently ignored for further investigation. A default choice may be $\alpha=0.05$; a smaller choice can be used if the tester is more risk-averse. With this, one can then plan the required debugging effort by testing all combinations with probabilities greater than $\alpha$ (call this set $\mathcal{S}$), beginning with the ones with highest probabilities. When $\mathcal{S}$ is small, this suggests just a small debugging effort can confidently identify the root cause. When $\mathcal{S}$ is large, this suggests a massive debugging effort may be needed, which indicates the software is not ready for release in the near future. Such insights can help test engineers forecast the required resources and timeframe for reliable software development.
@@
5:For each TF combination $(\mathbf{i},\mathbf{j})$, enumerate minimal covers for the failed cases in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, then compute its posterior root cause probability using Equations (10) and (15).
@@ 5 Case Studies
We now investigate the effectiveness of BayesFLo in two case studies: the first on our motivating application on the Traffic Alert and Collision Avoidance System, and the second on the Vulnerable Road User protection tests for safe autonomous driving.
@@ 5.1 Case Study 1: Traffic Alert and Collision Avoidance System
Consider first the motivating case study in Section 2, on the fault localization of the Traffic Alert and Collision Avoidance System (TCAS). TCAS targets the avoidance of potential mid-air collisions between aircraft, and its reliable performance is essential for mitigating catastrophic disasters (see, e.g., Davies and Yousif, 2025). In what follows, we consider a specific program – the Siemens test suite in Griesmayer et al. (2007) – for fault localization. Its $I=12$ input factors with corresponding levels are summarized in Table 1. We adopt the “TCAS v2” set-up from Griesmayer et al. (2007), which contains a fully-documented fault that results in an incorrect resolution advisory; if unaddressed, such a fault may lead to potential aircraft collisions. Here, it is known (Griesmayer et al., 2007) that such a failure arises from a single root cause, namely, the 3-FC involving factors 12, 8 and 9:
$\small(\texttt{Climb\_Inhibit}=1)\ \times\ (\texttt{Up\_Separation}=399)\ \times\ (\texttt{Down\_Separation}=640).$ (16)
This failure was found to arise as follows. Imagine two aircraft (one primary, the other incoming) are headed towards each other. Suppose the incoming aircraft is currently below but is projected to be above the primary aircraft at the closest approach (Bradley, 1992). When the climb rate is restricted for the primary aircraft (i.e., $\texttt{Climb\_Inhibit}=1$), there is a limit to the height that it can climb. In TCAS, a constant is thus added to the projected vertical separation between aircraft at closest approach due to an upward maneuver of the primary aircraft (i.e., Up_Separation), to ensure a resolution advisory is raised if it exceeds a projected downward maneuver (i.e., Down_Separation). Here, an incorrect constant was added in the system, resulting in resolution advisories not being raised when they should have been, thus precipitating a failure. We will investigate the effectiveness of BayesFLo compared to the state-of-the-art in identifying the root cause (16) with limited test runs.
With this set-up, we then perform test runs using the strength-2 covering array from Table 1, with a run size of $M=19$. Table 2 shows the corresponding test outcomes for each test run. Table 3 summarizes the ranked suspicious combinations using the BEN approach (Ghandehari et al., 2020) and the weighting approach (Lekivetz and Morgan, 2018), which serve as the state-of-the-art for combinatorial fault localization in our experiments. Recall the two limitations noted in Section 2. First, both approaches need to investigate eight suspicious combinations before identifying the true root cause, because they both require the inspection of 2-factor combinations before 3-factor ones. This can be undesirable given the costly nature of debugging each combination (Landsberg and Barr, 2018). Second, while both approaches provide a ranked list of suspicious combinations, the employed metrics cannot be interpreted as probabilities conditional on test data, which may hinder actionable decisions for confident fault localization. We explore next how BayesFLo can tackle these two limitations.
To implement BayesFLo, we need a careful specification of its prior distribution in (2). Here, prior domain knowledge on the nature of the incorrect advisory fault (described earlier) suggests that it is likely associated with the projected vertical separation between two aircraft, which involves the two “separation” factors Up_Separation and Down_Separation. To reflect this, we adopt the following informative prior. We set $p_{(i,j)}=1/30$ as the baseline prior probability, where its denominator $\sum_{i=1}^{I}J_{i}=30$ is the total number of levels for all $I=12$ factors. This baseline prior is assigned to all factors except for Up_Separation and Down_Separation; the latter is assigned a prior of $p_{(i,j)}=2/30$ to reflect heightened prior suspicion. We will contrast this with an “uninformative” implementation of BayesFLo, where all priors are assigned the same prior probability $p_{(i,j)}=1/30$. Posterior root cause probabilities are then computed via Algorithm 1. For the weighting approach in Lekivetz and Morgan (2018), the weights for the two “separation” factors are similarly set to be double that for the other factors to reflect this prior domain knowledge.
Combination Type Count Probability
3-FC 1 0.55
2-FC 1 0.41
2-FC 7 0.10
3-FC 140 <<0.01
(a) BayesFLo (uninformative prior).
Combination Type Count Probability
3-FC 1 0.65
2-FC 1 0.42
2-FC 2 0.14
2-FC 5 0.07
3-FC 140 <<0.03
(b) BayesFLo (informative prior).
@@
Table 4: Suspicious 2-FCs and 3-FCs ranked by their posterior root cause probabilities using BayesFLo (with uninformative and informative priors) for the TCAS case study. Combinations at the top are more suspicious and are investigated first. The true root cause is bolded.
Table 4 shows the 2-factor and 3-factor suspicious combinations from BayesFLo (using the above informative and uninformative prior specifications), ranked by their posterior root cause probabilities. Here, for all compared methods, we consider only suspicious combinations with at most three factors; this not only reduces computation, but is also guided by prior knowledge that root causes are typically of lower order for such systems (Kuhn et al., 2004). Out of 149 suspicious 2-FCs and 3-FCs, BayesFLo (with an informative prior) is able to pinpoint the true 3-factor root cause (16) as its most suspicious combination, with a posterior probability of 0.65. Two 2-FCs are ranked next with probabilities of 0.42 and 0.14, the remaining 2-FCs are ranked next with probabilities 0.07, and the remaining 3-FCs are ranked last with probabilities less than 0.03. From a debugging perspective, the BayesFLo analysis is preferable to existing techniques: it finds the true root cause as its top-ranked combination, without needing to further investigate and rule out the eight 2-FCs via expensive tests. This shows the importance of a Bayesian statistical framework that captures domain and structural knowledge, allowing for a principled comparison of suspicious combinations across different orders for efficient fault localization. Comparing BayesFLo with the informative vs. uninformative prior, we see that the incorporation of domain knowledge (i.e., the heightened suspiciousness of the two separation factors) raises the posterior probability of the true root cause, which shows that incorporating such knowledge can indeed enhance fault localization.
It is also important to investigate the probabilistic aspect of the BayesFLo analysis, and how this can help guide actionable decisions for confident fault localization. Recall that for existing analyses (see Table 3), which do not employ probabilistic metrics, it may be difficult to plan the subsequent debugging effort, since it is not clear which 2-FCs or 3-FCs can be confidently ignored for debugging. Here, BayesFLo can give a clearer picture. Using the recommended significance level $\alpha=0.05$, the (informative) BayesFLo analysis suggests only nine combinations need to be tested; the last 140 3-FCs can be confidently ignored, since their posterior probabilities are less than $\alpha$. This is quite reasonable given test outcomes, where only two failed test runs were observed. The probabilistic analysis from BayesFLo thus enables effective planning of debugging efforts for confident fault localization.
5.2 Case Study 2: Vulnerable Road User Protection Tests
Consider next a case study on the Vulnerable Road User protection test for safe autonomous driving. This case study is adapted from Gladisch et al. (2020), where the goal is to test and validate safety standards for automated driving systems via the Vulnerable Road User protection tests. VRU tests evaluate how effective the autonomous emergency braking system is at detecting and reacting to pedestrians and cyclists under a broad range of driving conditions. The VRU scenario considered in Gladisch et al. (2020) involves a child jaywalking across an intersection in front of an approaching vehicle. Here, we consider a system with $I=15$ factors, listed in Table 5 with their corresponding levels. The first ten factors were investigated in Gladisch et al. (2020), and the last five are new factors that we added in our case study. Unfortunately, Gladisch et al. (2020) explores the problem of test case design and does not shed light on specific root causes. As a proof-of-concept, we consider the following 2-FC as the root cause:
$(\texttt{Reflection on road}=\texttt{yes})\ \times\ (\texttt{Speed}=\texttt{fast}).$ (17)
This is quite intuitive: under fast driving speeds and high reflection from the road surface, the automatic driving system can fail to detect the jaywalking child, resulting in a failure.